OpenStudy (anonymous):

what is the quotient of 5/18 divided by 5/12

OpenStudy (owlcoffee):

Having this: \[\frac{ \frac{ 5 }{ 18 } }{ \frac{ 5 }{ 12 } }\] or for more gentle looking: \[\frac{ 5 }{ 18 }\div \frac{ 5 }{ 12 }\] Now there's a little trick in algebra, for fractionary divisions, let's take a look at it: \[\frac{ a }{ b } \div \frac{ c }{ d } = \frac{ a }{ b }\times \frac{ d }{ c }\] Now, what does it mean? Well, it means I cant transform the division sign to a multiplication sign, if I switch places on the second fractions... Let's try that: \[\frac{ 5 }{ 18 }\div \frac{ 5 }{ 12 }=\frac{ 5 }{ 18 }\times \frac{ 12 }{ 5 }\] Now, let's take a look at the fractionary multiplication: \[\frac{ a }{ b }\times \frac{ c }{ d }=\frac{ ac }{ bd }\] It only means that you multiply the numbers above, and the numbers below all together. Let's apply that: \[\frac{ 5 }{ 18 }\times \frac{ 12 }{ 5 }=\frac{ (5)(12) }{ (18)(5) }\] All we have to do now is multiply :) \[\frac{ (5)(12) }{ (18)(5) }=\frac{ 60 }{ 90 }\] And there we got it :) That fractions can be simplified yes, but I'll leave that to you.